JAVA What is the running time required for repeatedly splitting a problem into two equally sized halves, assuming no additional work needs to be done?

O(1)

O(log10N)

O(log2N)

O(log2N)

O(N)

O(N log N)

O(N2)

O(N3)

O(Nk)

O(2N)

O(N!)
Answer #1
Recursive formula for repeatedly splitting a problem into two equally sized halves, assuming no additional work needs to be done
T(n) = T(n/2) + 1
     = T(n/4) + 1 + 1
     = T(n/8) + 1 + 1 + 1
     ......
     ......
     ......
     = T(n/n) + 1 + .... + 1 + 1 + 1 [log(n) +1 terms]
     = 1 + 1 + .... + 1 + 1 + 1 [log(n) +1 terms]
     = log2(n)
     = O(log2n)

O(log2N)